Starting from rest a 0.064-kg arrow is fired horizontally. The bowstring exerts a horizontal average force of 81 N on the arrow over a distance of 0.77 m. With what horizontal speed does the arrow leave the bow?
1
Expert's answer
2020-05-21T13:59:29-0400

From the conservation of energy:


"0.5mv^2=Fs"

"0.5(0.064)v^2=(81)(0.77)"

"v=44\frac{m}{s}"